Can someone help me? im stuck on this...
Write three quadratic equations, with a, b, and c (coefficients of x2, x, and the constant) as:
1...Integers
2...Rational numbers
3....Irrational numbers

Thank you!

Answer by Alan3354(69443)

Write three quadratic equations, with a, b, and c (coefficients of x2, x, and the constant) as:
1...Integers
2x^2 + 3x + 4 = 0
----------------
2...Rational numbers
(2/3)x^2 - (1/7)x + 1/9 = 0
---------------
3....Irrational numbers
pi*x^2 + e*x + sin(33º) = 0